# IntersectionalFairness.bib

### v1

@@ 1 Introduction
Our contribution is fourfold: We (1) elucidate the problem of intersectionality in algorithmic fairness, (2) develop desiderata to clarify the challenges that underlie the problem of intersectionality and to guide the search for potential solutions. Moreover, we (3) illustrate the desiderata and potential solutions by presenting a statistical setup that uses simple hypothesis testing, and (4) evaluate this proposal, partly empirically, in light of the desiderata.
Our topic is an under-appreciated problem in non-ideal epistemology [mckenna_non-ideal_2023]. However, our larger aim is to advance the literature on algorithmic fairness more broadly. The approach that we propose in response to the problem of intersectionality differs fundamentally from the typical way of “measuring” algorithmic fairness. [^1] We use fairness “measure,” “metric” and their cognates with two caveats. First, the problem is one of estimation not measurement. Second, fairness metrics are meta-metrics since they aggregate a higher-dimensional vector of model performance into a lower-dimensional summary [lum_-biasing_2022]. We hence advance the debate, by pointing out possibilities of approaching fairness differently: as accounting for uncertainty (instead of concentrating on point estimates) and as a matter of sufficiency (instead of equality).
@@ 2.2 The Problem of Intersectionality
As the number of attributes that define subgroups grows, the amount of data available for each subgroup shrinks rapidly. After all, the number of subgroups grows exponentially with the number of protected attributes: For $n$ binary attributes, there are $2^{n}$ intersectional groups. This, in turn, entails a data problem: When social identities are constituted by intersections of increasingly many variables, and when the constituting attributes are not just binary, the data within each of the intersections can become very small. In Europe, where discrimination is highly intersectional and fairness audits are encouraged by legislation, [^2] Recital 49 of the EU Artificial Intelligence Act [EUAI] encourages “the development of benchmarks and measurement methodologies for AI systems” [Recital49]. Yet the statistical problems of intersectional fairness are, in some way, greater in Europe. Since nationality groups are already comparatively small, intersectional groups are even smaller subgroups within already small nationality groups. For example, Hungarian Roma face discrimination in the housing market, Maghrebi French in the labor market, whereas people of African descent in England and Wales face discrimination in the criminal justice system [IntersectionalityEurope]. fairness audits may need to account for several thousand subgroups. [^3] Assuming 3 binary attributes (e.g., non-white, cis-gender, same-sex orientation) 1 three-valued attribute (e.g., gender as ‘male,’ ‘female,’ and ‘neither’), 9 different ethnic backgrounds (e.g., Roma, Chinese, Turkish), and 12 different nationalities or localities (e.g., Hungarian, German, French), yields 2,592 intersectional subgroups. And this number may be conservative since the number of discernible ethnic groups is larger than 9, of nationalities is larger than 12, and the legally protected attribute of age is not even included. And because gathering the data necessary for fairness audits is typically costly— e.g., the “ground truth” needs to be established to assess whether a prediction is correct—such data tend to be scarce to begin with.
@@ 3 Existing Work
Essentially, the addition of $\alpha(G)$ relaxes the original fairness metric of (1) depending on the proportion of $G$ as a share of the overall population. The smaller $G$ is, the more the condition is relaxed. As Kearns et al. explain, this addition is necessary to enable statistical estimation, given the increasing statistical uncertainty with decreasing group size. We discuss the implications in Section 4.
@@ 4 Desiderata
Although the problem of intersectionality is recognized in the literature, how difficult this problem is may not have been fully appreciated. At least some of the existing contributions violate minimal moral or methodological desiderata, as we shall see in Sections 4.1, 4.2, and 4.3.
@@ 4.2 Consistent Conceptualization
This desideratum may resemble that of Minimal Justice. But whereas Minimal Justice is a moral desideratum, Consistent Conceptualization is a methodological one. Consistent Conceptualization aims to ensure minimal construct validity, i.e., that a certain metric represents a concept of interest correctly. The importance of construct validity for fairness is already established in the literature [jacobs_measurement_2021].
@@ 4.3 Incentive Compatibility
This is a “perverse” effect because, in practice, additional data collection about a minority group will help improve the model performance for that group. In other words, the metric gives an incentive to do the opposite of what it is meant to achieve.
Whether other metrics [AuditingAchievingIntersectional, 9101635, MolinaLoiseauBoundingApproximating] violate this desideratum depends on whether the estimated performance disparity is greater than the true disparity (which further data would likely help approximate). Fairness metrics that operationalize fairness as equality (e.g. as model performance disparity across groups), incentivize $m(G)$ to be nearly the same for all subgroups $G$. If the true model performance is nearly equal among groups, then these metrics incentive further data collection in order to have more accurate estimates of $m(G)$.
@@ 5.2 Pessimist’s Model
Depending on a model’s deployment context, the optimistic approach might be problematic. Consider instead the following pessimistic hypothesis test.
@@ 6 Fairness Datasets Analysis
To evaluate whether our metrics in practice incentivize or disincentivize additional data collection, we assess their results on well-known “fairness datasets.” To do this, we use lale; a Python library created by IBM [lale_library]. This library allows for creation of consistent automated machine learning models. It also has 20 fairness datasets that can easily be fetched, modeled, and evaluated [lale_fairness_datasets]. These 20 datasets are all tabular with a categorical target variable. They each come with “fairness metadata,” which includes protected attributes, along with ranges/values of those attributes that correspond to the privileged group. [^13] While there have been critiques of the usage of some of these datasets [ding2021retiring] [compaslicated], they are still appropriate for the purpose of testing whether our proposals incentivize or disincentivize the collection of additional data.
@@
For present purposes, to concentrate on the philosophical contributions of the project, we leave details on the methods of our analysis to Appendix C and only highlight the main result. Figure 1 shows the results of subsampling only the critical subgroup, that is the one that is on the $c$ threshold. [^14] Some of the plots are disconnected. This is because sometimes the subsampling of the dataset did not include any members of the critical subgroup; in those cases, the model could not predict for that subgroup, so no accuracy measurement could be taken. The most erratic curves (curves of $m(G)$ and $c_{2}$ for the creditg and nlsy datasets) correspond to either a subgroup of size 1 or 2.
@@
Figure 1: Plots of $m(G),c_{1}^{g}$, and $c_{2}^{g}$ of critical subgroups $G$ for each dataset. Here we subsampled only the critical subgroup, and the $x$-axis corresponds to the percentage of the critical subgroup that is kept. Legend lists the dataset name.
@@
The thing to observe here is that there is a trend upwards in each of these plots. This suggests that our optimist’s metric—at least for the datasets tested—does not pose perverse incentives. The further we go on the x-axis (representing more data being “collected”), the model performance as well as the fairness metrics tend to improve.

### v2

@@ 1 Introduction
Our contribution is fourfold: We (1) elucidate the problem of intersectionality in algorithmic fairness, and (2) develop desiderata to clarify the challenges that underlie the problem of intersectionality and to guide the search for potential solutions. Moreover, we (3) illustrate the desiderata and potential solutions by presenting a statistical setup that uses simple hypothesis testing, and (4) evaluate this proposal, partly empirically, in light of the desiderata.
Our larger aim is to advance the literature on algorithmic fairness more broadly. The approach that we propose in response to the problem of intersectionality differs fundamentally from the typical way of “measuring” algorithmic fairness. [^1] We use fairness “measure,” “metric” and their cognates with two caveats. First, the problem is one of estimation not measurement. Second, fairness metrics are meta-metrics since they aggregate a higher-dimensional vector of model performance into a lower-dimensional summary [lum_-biasing_2022]. We hence advance the debate, by pointing out possibilities of approaching fairness differently: as accounting for uncertainty (instead of concentrating on point estimates) and as a matter of sufficiency (instead of equality).
@@ 2.2 The Problem of Intersectionality
As the number of attributes that define subgroups grows, the amount of data available for each subgroup shrinks rapidly. After all, the number of subgroups grows exponentially with the number of protected attributes: For $n$ binary attributes, there are $2^{n}$ intersectional groups. This, in turn, entails a data problem: When social identities are constituted by intersections of increasingly many attributes, and when these constituting attributes are not just binary, the data within each of the intersections can become very small. In Europe, where discrimination is highly intersectional and fairness audits are encouraged by legislation, [^2] Recital 49 of the EU Artificial Intelligence Act [EUAI] encourages “the development of benchmarks and measurement methodologies for AI systems” [Recital49]. Yet the statistical problems of intersectional fairness are, in some way, greater in Europe. Since nationality groups are already comparatively small, intersectional groups are even smaller subgroups within already small nationality groups. For example, Hungarian Roma face discrimination in the housing market, Maghrebi French in the labor market, whereas people of African descent in England and Wales face discrimination in the criminal justice system [IntersectionalityEurope]. fairness audits may need to account for several thousand subgroups. [^3] Assuming 3 binary attributes (e.g., non-white, cis-gender, same-sex orientation) 1 three-valued attribute (e.g., gender as ‘male,’ ‘female,’ and ‘neither’), 9 different ethnic backgrounds (e.g., Roma, Chinese, Turkish), and 12 different nationalities or localities (e.g., Hungarian, German, French), yields 2,592 intersectional subgroups. And this number may be conservative since the number of discernible ethnic groups is larger than 9, of nationalities is larger than 12, and the legally protected attribute of age is not even included. And because gathering the data necessary for fairness audits is typically costly—e.g., the “ground truth” needs to be established to assess whether a prediction is correct—such data tend to be scarce to begin with.
@@
Our discussion hence adds to the existing technical and critical objections against (intersectional) algorithmic fairness [corbett-davies_measure_2024, kong_are_2022], acknowledging that a commitment to intersectionality and fairness likely requires a broader set of actions than estimating certain properties of models [wang_towards_2022, suresh_towards_2022, klumbyte_critical_2022].
@@ 3 Existing Work
Essentially, the addition of $\alpha(G)$ relaxes the original fairness metric of (1) depending on the proportion of $G$ as a share of the overall population. The smaller $G$ is, the more the condition is relaxed. As Kearns et al. explain, this addition is necessary to enable statistical estimation, given the increasing statistical uncertainty with decreasing group size. We discuss the implications in Section 4, and give the results of an empirical study regarding this formulation in Appendix E.
@@ 4 Desiderata
Although the problem of intersectionality is recognized in the literature, how difficult this problem is may not have been fully appreciated. At least some of the existing contributions violate minimal moral or methodological desiderata, as we shall see in Sections 4.1, 4.2, and 4.3 (and Appendix E).
@@ 4.2 Consistent Conceptualization
This desideratum may resemble that of Minimal Justice. But whereas Minimal Justice is a moral desideratum, Consistent Conceptualization is a methodological one. Minimal Justice assumes a standard of fairness as given (and requires that it not be lower for certain groups). Consistent Conceptualization ensures that this standard has minimal construct validity, i.e., that the formal operationalization of fairness represents the informal, intended conception of fairness (whichever that may be). The importance of construct validity for fairness is already established in the literature [jacobs_measurement_2021].
@@ 4.3 Incentive Compatibility
This is a “perverse” effect because, in practice, additional data collection about a minority group will help improve the model performance for that group. In other words, the metric gives an incentive to do the opposite of what it is meant to achieve. [^10] We do not contend that more data should be collected. Privacy considerations are important. Our point is instead that maintaining the appearance of a good fairness metric is a bad reason to not collect more data.
Whether other metrics [AuditingAchievingIntersectional, 9101635, MolinaLoiseauBoundingApproximating] violate this desideratum depends on whether the estimated performance disparity is greater than the true disparity (which further data would likely help approximate). Fairness metrics that operationalize fairness as equality (e.g., as model performance disparity across groups), incentivize $m(G)$ to be nearly the same for all subgroups $G$. If the true model performance is nearly equal among groups, then these metrics incentive further data collection in order to have more accurate estimates of $m(G)$.
@@ 5.2 Pessimist’s Model
Depending on a model’s deployment context, the optimistic approach might be problematic. [^13] Depending on the ethical risks involved in how a model is used, the more precautionary assumptions behind the pessimist’s model might be more appropriate. Consider instead the following pessimistic hypothesis test.
@@ 6 Fairness Datasets Analysis
We evaluate empirically whether our metrics meet the desideratum of incentive compatibility. The question is: Do our metrics incentivize or disincentivize additional data collection?
To answer this question, we “simulate” additional data collection by experiment. We train models on increasingly larger subsamples of benchmark datasets and observe how metrics behave as the size of the training data increases. The behavior that we want to see is that the fairness metrics increase with the size of the training data sampled from the dataset. If, instead, a fairness metric decreased as greater shares of the dataset are sampled, the metric would disincentivize further data collection.
@@
We seek to observe our metrics’ behavior across the largest feasible range of benchmarks. To achieve this, we use lale, a Python library created by IBM [lale_library]. Lale allows for the creation of consistent automated machine learning models across 20 well-known “fairness datasets” that can easily be fetched, modeled, and evaluated [lale_fairness_datasets]. These datasets are all tabular with a categorical target variable. They each come with “fairness metadata,” which includes protected attributes, along with ranges/values of those attributes that correspond to the privileged group. [^15] While there have been critiques of the usage of some of these datasets [ding2021retiring] [compaslicated], they are still appropriate for the purpose of testing whether our proposals incentivize or disincentivize the collection of additional data. Details on the methods of our analysis are in Appendix C. Here we only discuss the main result on testing whether our metrics incentivize against data collection.
For each of the datasets, we observe model performance $m(G)$, as well the optimist’s $c_{1}^{g}$ and the pessimist’s fairness metric $c_{2}^{g}$ respectively. For ease of interpretation we use accuracy as model performance; neither our results nor their interpretation depend on this.
@@
We ran two versions of this experiment. In one version, we subsample the entire dataset of each benchmark; whereas in another, we subsample only on the critical subgroup, which is the group that is right on the $c$ threshold. The first version simulates additional data collection for all groups, whereas the latter for those groups that “drag down” the fairness metric. Here we concentrate on results from subsampling on the critical subgroup only, shown in Figure 1.161616Some of the plots are disconnected. This is because sometimes the subsampling of the dataset did not include any members of the critical subgroup; in those cases, the model could not predict for that subgroup, so no accuracy measurement could be taken. The most erratic curves (curves of $m(G)$ and $c_{2}$ for the creditg and nlsy datasets) correspond to either a subgroup of size 1 or 2. Full results for both versions are in Appendix C.4.
@@
Figure 1: Plots of accuracy $m(G)$, optimist’s metric $c_{1}^{g}$, and pessimist’s metric $c_{2}^{g}$ of critical subgroups $G$ for each dataset. The $x$-axis corresponds to the percentage of the critical subgroup that is kept. Legend lists the dataset name.
@@
The thing to note here is that there is a trend upwards in each of these plots. Most notably, the middle plot, on the optimist’s metric $c_{1}^{g}$ shows this upward trend. [^17] Some datapoints “overshoot” on the $y$-axis with values $>1$, suggesting a negative trend, e.g., for the compas dataset (green line). But this behavior is an artifact of the standard way of calculating the confidence interval. This suggests that our optimist’s metric—at least for the datasets tested—does not pose perverse incentives. The further we go on the x-axis (representing more data being “collected”), the model performance as well as the fairness metrics tend to improve.
Consider for example the behavior of the optimist’s metric for the model trained on increasing amounts of data from the tae dataset (brown line that “starts” lowest in middle figure). Although the metric does not strictly increase as the training is based on greater data (the metric decreases slightly from 20% to 30% of data used), it shows a very strong upward trend.
@@ 7 Conclusion
Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. DMS-1928930 and by the Alfred P. Sloan Foundation under grant G-2021-16778, while Ellen Veomett was in residence at the Simons Laufer Mathematical Sciences Institute (formerly MSRI) in Berkeley, California, during the Fall 2023 semester.
@@ Appendix D Tables
Appendix E Analysis of Metric from Kearns et al.
As noted in Section 4.3, we hypothesize that the fairness metric outlined by Kearns et al. [Kearns_FairnessGerrymandering] violates Incentive Compatibility. The fairness metric likely “looks worse” as additional data is gathered about a small subgroup (i.e., a group whose size in proportion to the entire dataset is small). The fairness metric includes a factor which is the proportion of the subgroup within the dataset. Thus, as additional data is collected from that subgroup alone, this proportion increases, making the model more likely to violate the fairness criteria, hence potentially disincentivizing additional data collection on that subgroup. Here we empirically examine this hypothesis.
E.1 Study Description
Using the same datasets, subgroups, pre-processing, cleaning, and models outlined in Appendix C, we calculate the value of the following expression from Equation (2):
$\alpha(G)|m(G)-m(\cdot)|$ (6)
Recall that $\alpha(G)$ is the proportion of group $G$ within the total population, and that $m$ is some model performance metric (as in Appendix C, we use accuracy as our sample metric $m$ for this study). The value $m(G)$ is the model performance metric evaluated only on subgroup $G$, while $m(\cdot)$ is the value of the model performance metric on the entire dataset.
Kearns et al. [Kearns_FairnessGerrymandering] use an auditing process wherein the value calculated from expression (6) must be below some threshold $\epsilon$ in order for a model to be considered fair. Thus, we can think of expression (6) as describing unfairness for group $G$ [^20] All typical ways of defining “fairness” can be interpreted this way. A higher $\epsilon$ in (1) is interpreted as a decrease in fairness and thus an increase in unfairness.
E.2 Methods
We calculate the value in expression (6) on increasing subsamples of each dataset. We concentrate on small sugbroups $G$ that comprise no more than 10% of the total population. Just as in the experiments described in Appendix C, we examine two subsampling scenarios: We subsample the subgroup in question only (to simulate gathering more subgroup data), and we subsample the entire dataset (to simulate gathering more population data). Note that, of course, the values of $m$ depend on the model created, which depends on the subsample of the data used to create the model.
Many of the datasets (heart_disease, nursery, ricci, student_math, student_por, tae and us_crime) don’t have any subgroups comprising less than 10% of the total population, and thus we exclude those datasets from this analysis. From the other datasets, we concentrate on four: the adult, bank, meps20, and titanic datasets. The results for all other datasets are similar.
In Section 4.3, we hypothesize that the protocol of Kearns et al. violates Incentive Compatibility. Specifically, we hypothesize that when subsampling only small groups, the unfairness value of expression (6) would increase. Since the value $\alpha(G)$ does not change significantly when subsampling the entire population, we do not expect expression (6) to change much when subsampling the entire dataset, aside from the fact that potentially a better model might make expression (6) decrease.
E.3 Results and Discussion
The results of subsampling just the small subgroup can be found in Figure 4, and the results of subsampling the entire dataset can be found in Figure 5.
When only the small subgroup is subsampled (as in Figure 4), we see the value of (6) increasing for all subgroups in the adult and bank datasets. The picture is slightly more muddled in the meps20 and titanic datasets, but these still show either a consistent increase or an initial increase for nearly all of the small subgroups. In other words, the value (6) of unfairness increases, indicating that the Kearns et al. auditing process discourages additional data collection of small subgroups, and thus violates Incentive Compatibility.
Figure 4: Values of expression (6) on the adult, meps20, titanic, and bank datasets. Horizontal axis is the percent of the subgroup, vertical axis is unfairness (i.e., the value of expression (6)).
When the entire dataset is subsampled (as in Figure 5), values of (6) remain remarkably consistent in the adult dataset, and tend to decrease in the bank, meps20, and titanic datasets. We can thus conclude that the Kearns et al. approach, while it discourages collecting additional data from only the smallest subgroups in a dataset (thereby not satisfying Incentive Compatibility), does not appear to discourage additional data collection when each subgroup’s proportion within the population stays consistent.
Figure 5: Values of expression (6) on the adult, meps20, titanic, and bank datasets. Horizontal axis is the percent of the entire dataset kept, vertical axis is unfairness (i.e., the value of expression (6)).